OpenStudy (anonymous):

Use the matrix tool to solve the system of equations. Choose the correct ordered pair. 3x + 7y = 1 -x + y = -7 A. (4, -3) B. (3, -4) C. (2, -5) D. (5, -2)

Parth (parthkohli):

-3x + 3y = -21 Adding equations. 10y = -20 y = -2 -x + -2 = -7 =-5 = -x So, x = 5

Parth (parthkohli):

D. 5, -2

Parth (parthkohli):

Sorry that I didn't use the matrix tool. I dunno that meth.